OpenStudy (anonymous):

a parabola has a vertex of (-5,0) and passes through point (-3,1) which of the following is the equation of this parabola in the form y=a(x-h)^2+k

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

h=-5, a=1/4 and k=0 so: y=1/4(x+5)